### v1

# Credit Risk Meets Large Language Models: Building a Risk Indicator from Loan Descriptions in P2P Lending

Abstract.
Peer-to-peer (P2P) lending has emerged as a distinctive financing mechanism, linking borrowers with lenders through online platforms. However, P2P lending faces the challenge of information asymmetry, as lenders often lack sufficient data to assess the creditworthiness of borrowers. This paper proposes a novel approach to address this issue by leveraging the textual descriptions provided by borrowers during the loan application process. Our methodology involves processing these textual descriptions using a Large Language Model (LLM), a powerful tool capable of discerning patterns and semantics within the text. Transfer learning is applied to adapt the LLM to the specific task at hand.
Our results derived from the analysis of the Lending Club dataset show that the risk score generated by BERT, a widely used LLM, significantly improves the performance of credit risk classifiers. However, the inherent opacity of LLM-based systems, coupled with uncertainties about potential biases, underscores critical considerations for regulatory frameworks and engenders trust-related concerns among end-users, opening new avenues for future research in the dynamic landscape of P2P lending and artificial intelligence.
Credit risk, P2P Lending, BERT, Transfer learning, Explainable AI, XGBoost
journal: TMISccs: Computing methodologies Information extractionccs: Applied computing Business process management systems
1. Introduction
Peer-to-peer (P2P) lending is a growing phenomenon that allows individuals to engage in direct lending and borrowing transactions, bypassing traditional financial institutions. The process is facilitated through online platforms, where prospective borrowers submit loan applications and potential lenders make informed decisions about where to invest their funds.
An inherent challenge in P2P lending is the presence of information asymmetry, wherein borrowers possess more and often superior information compared to lenders. To address this issue, platforms employ strategies to complement the conventional data provided in loan applications (Cummins et al., 2019). For instance, borrowers are frequently encouraged to provide a voluntary textual description describing the purpose of the loan and their particular situation. Despite the absence of formal verification, such voluntary disclosures have been observed to stimulate increased bidding activity among lenders. However, lenders may lack the expertise to effectively assess the creditworthiness of borrowers and may be influenced by different factors (Michels, 2012).
Traditional credit scoring models do not harness the rich information embedded in the narratives submitted by loan applicants. Several approaches have tried to incorporate such information into credit risk models. The approaches include the extraction of linguistic metrics (Gao and Lin, 2015; Dorfleitner et al., 2016), the use of topic modeling to characterize the underlying themes (Xia et al., 2020; Zhang et al., 2020), or a combination of both (Siering, 2023).
In our work, we aim to use state-of-the-art Large Language Models (LLMs) to analyze textual descriptions and assign them a risk score. To achieve this, we will use BERT (Bidirectional Encoder Representations from Transformers) (Devlin et al., 2018), renowned for its outstanding capabilities in comprehending contextual nuances and semantic intricacies within textual data. BERT has demonstrated success in fine-tuning for classification tasks (Sun et al., 2019) and adapting to specific domains, such as biomedicine (Lee et al., 2019), as well as specialized tasks like spam classification (Tida and Hsu, 2022).
We will apply transfer learning to train BERT in distinguishing between default and non-default loans based on their textual descriptions. Subsequently, we will explore the viability of using the BERT output as a risk score and analyze how it relates to other variables. We integrate the BERT score into a credit risk classifier and show that it improves its predictive performance. Our results open new avenues for research by emphasizing the potential of LLMs to enhance credit risk models in P2P lending. Moreover, they underscore the need for greater transparency to streamline the adoption of these approaches, fostering trust among users and regulatory entities.
This paper is organized as follows: Section 2 provides a comprehensive review of related work in the field of credit risk assessment and natural language processing. Section 3 presents an overview of LLMs and the BERT model. Section 4 describes the dataset used, detailing the data preprocessing steps and conducting an in-depth exploratory data analysis. Section 5 outlines the methodology, model architecture, and training procedures employed in integrating BERT into the credit risk assessment framework. Section 6 provides an analysis of the risk score generated by the BERT description processing. Section 7 discusses the results of our experiments, highlighting the improvements achieved by incorporating BERT-based textual analysis. Finally, Section 8 concludes the paper by summarizing key findings, discussing implications, and suggesting avenues for future research in the intersection of NLP and credit risk assessment.
2. Related work
In their comprehensive analysis of risk-return modeling within the P2P lending market (Ariza-Garzón et al., 2021), the authors identify a discernible trend towards including new sources and types of information to improve risk and profit management models in the P2P market. The sources are very diverse and include transactional data (Xu et al., 2022), the topology of the lending-borrowing network (Li et al., 2018), data from social networks (Xu et al., 2016), or, more recently, facial features (Qi et al., 2022). Among them, the authors identify as a predominant trend the inclusion of textual data taken from statements describing the purpose of the loan.
In a pioneer work exploring the impact of textual factors on peer-to-peer lending (Herzenstein et al., 2011), the authors analyze P2P loans including manually annotated narrative aspects, such as trustworthiness, economic hardship, hard work, success, morality, and religiosity. These aspects were combined with demographic variables and loan characteristics. Their results highlight that narratives regarding trustworthiness have a significant influence on decision-makers, i.e. credit lenders, in terms of securing loan funding, and that some of these narratives also play a significant role in loan performance.
However, most subsequent studies typically use text mining or artificial intelligence methods to extract linguistic features or loan description topics. Regarding the use of linguistic features, the authors in (Gao and Lin, 2015) use machine learning and text mining techniques to quantify and extract linguistic features (e.g., readability, positivity, objectivity, and deception cues), and then build both explanatory econometric models and predictive models using such features. They find that they can indeed reflect borrowers’ creditworthiness and predict loan default. They also use a panel of investors and confirm that investors indeed value texts written by borrowers, but that they can also be deceived by some of the deception cues well established in the literature. Similarly, in (Dorfleitner et al., 2016), the authors include linguistic factors and the presence of social and emotional keywords and evaluate their impact on two European platforms. They found that text-derived variables influence the probability of funding, but not the probability of default. In (Wang et al., 2016), linguistic statistical features and abstract text features (including deception, subjectivity, sentiment, readability, personality, and mindset) are used to characterize text descriptions. They compare the performance of different classifiers based on the textual features and conclude that their performance is close to that of the classifiers using traditional financial features, but that adding textual features can improve the performance of the whole credit risk evaluation system.
In topic modeling, the Latent Dirichlet Allocation model (LDA) has been widely used. In (Jiang et al., 2017), the authors use LDA to extract six topics from the loan descriptions whose meanings are obvious: assets, income and expenses, work, family, business, and agriculture. They also consider the number of characters in the descriptive text. They conclude that soft (qualitative) information can improve the performance of loan default prediction compared to existing methods based only on hard (quantitative) information and that soft features have a significant ability to discriminate loan defaults. Similarly, in (Yao et al., 2018) an LDA topic model is used to classify the loan titles into six purposes. The results show that the purpose significantly affects a borrower’s likelihood of obtaining financing. In particular, ambiguous titles, where a borrower does not adequately explain the purpose of the loan, significantly reduce the likelihood of successfully securing the loan. In (Xia et al., 2020), Xia et al. used a keyword clustering algorithm to extract the topics automatically. The algorithm uses keyword extraction based on term frequency-inverse document frequency (TF-IDF) and word embeddings using the Word2Vec model (Mikolov et al., 2013) (a neural network). Their results on three real-world datasets show that topic variables significantly improve predictability compared to using traditional information alone.
More recently, Siering (Siering, 2023) investigated the effect of both aspects: topics and linguistic features. To extract topics, the author utilized a financial text analysis procedure (Loughran and McDonald, 2011) to build a domain dictionary. The identified topics describe the purpose of the loan, as well as the borrowers’ request for help, reliability, or appreciation. These topics were operationalized as binary indicator variables. The author employs text mining to create variables that measure aspects such as polarity, active orientation, readability, average sentence length, and word count. These variables are then used as input in a logistic regression. The results indicate that both linguistic and content-based factors aid in predicting a loan’s probability of default, with content-based factors being more significant than linguistic variables. The author also analyzes how the variables contribute to increasing or reducing the probability of default, and finds, for example, that expressing reliability has a positive effect.
Yet, a significant gap becomes apparent in the application of state-of-the-art natural language techniques, including deep learning methods. In their work (Zhang et al., 2020), Zhang et al. study the transformer encoder’s ability to capture textual features from loan descriptions. These characteristics, alongside the traditional hard features derived from loan applications, are inputted into a neural network to predict the likelihood of loan default. The study highlights the effectiveness of transformers, showing better results when including textual loan descriptions compared to models that do not consider them.
In this work, we propose that the potential improvement could be further enhanced by using encoder-based LLMs, such as BERT, which has been successful in improving classification in other fields.
3. An overview of Large Language Models and BERT
Large Language Models (LLMs) are built upon the Transformer architecture (Vaswani et al., 2017), leveraging attention mechanisms to enhance language comprehension. LLMs can be broadly categorized into three primary families, each distinguished by its architecture:
  • Encoder-only, widely employed for language comprehension tasks such as text classification, named entity recognition, and extractive question answering. The most famous example is BERT (Devlin et al., 2018), and we will explain it in more detail below.
  • Decoder-only, designed for generative tasks, exemplified by the well-known GPT models (Radford et al., 2019; Brown et al., 2020). It is employed in various tasks, including question answering (Pride et al., 2023), text summarization (Bhaskar et al., 2022), and programming code generation (Chen et al., 2021).
  • Encoder-decoder models, suited for tasks demanding both language understanding and generation, such as language translation or text summarization. The most influential models are BART (Lewis et al., 2019) and T5 (Raffel et al., 2020).
The selection of the appropriate architecture hinges on the specific requirements of the intended task. Whether it be the nuanced comprehension of language, creative text generation, or the synthesis of both, the versatility of LLMs offers a tailored solution for diverse applications.
We will focus on BERT (Bidirectional Encoder Representations from Transformer) which is a Transformer-based language model introduced by Google researchers in 2018 (Devlin et al., 2018). Its architecture consists of a stack of encoders from the Transformer model. The bidirectional nature of BERT is key, as it considers both the left and right context of each word, enhancing its ability to understand context-dependent meanings and to be effective in language understanding tasks. Numerous studies have consistently shown that BERT is the most effective linguistic model for various of these tasks (Kriebel and Stitz, 2022; Stevenson et al., 2021). In contrast, the widely recognized GPT (Radford et al., 2019; Brown et al., 2020) model is built with a Transformer decoder stack, making it more suitable for language generation tasks.
BERT stands as a milestone whose success has spurred the development of a diverse family of models that build upon its architecture. Some versions aim to achieve similar performance while having a smaller number of parameters, such as DistilBERT (a distilled version of BERT) (Sanh et al., 2019), or ALBERT (A Little BERT) (Lan et al., 2019). Others are adaptations to other languages such as CamemBERT (Martin et al., 2020) to French or BETO to Spanish (Cañete et al., 2020). Other proposals aimed to improve upon BERT by modifying some design decisions when pretraining BERT and also training the model longer, as in the case of RoBERTa (Robustly optimized BERT approach) (Liu et al., 2019), which resulted in improved contextualized representations and enhanced language understanding.
BERT has also been fine-tuned for specific tasks using transfer learning. For example, BioBERT (Lee et al., 2019) is a BERT model trained on biomedical documents to perform text mining tasks such as named entity recognition, relation extraction, and question answering. Other works fine-tune BERT to excel on text classification in well-known datasets (Sun et al., 2019) and adapted to specific applications on sentiment analysis (Gao et al., 2019) or spam classification (Tida and Hsu, 2022).
In our research, we aim to develop a credit risk model by leveraging a BERT-based model. Specifically, our model will utilize the textual descriptions of loans provided by borrowers in their application forms, producing a predictive score indicative of the likelihood of loan default. We will assess the effectiveness of the score in a granting model using the well-established dataset from the Lending Club P2P lending platform.
4. Description of the dataset
We have used a public data set of the P2P lending company Lending Club [^1] https://www.kaggle.com/wordsforthewise/lending-club, which is widely used in credit risk publications and the most widely used when dealing with the P2P market (Ariza-Garzón et al., 2021).
The dataset includes loans granted by the company between 2007 and 2018, with a total number of 2,260,699 instances. However, we only consider loans whose final status is known, that is, that were either fully repaid or defaulted. Thus, we eliminate loans in transitory states (in a grace period, late, etc.) and exclude loans with no information on income and indebtedness. This led us to a dataset containing 1,347,681 instances.
Furthermore, we only use the loans that include the text description (desc variable), which is crucial for our study. For unknown reasons, we noticed that the descriptions were no longer available as of March 2014, so we further filtered the dataset considering only the loans from the beginning (June 2007) until that month.
After the desc variable cleaning, our dataset consists of 119,101 obligations, all of them including a description and with a target binary variable called Default. The target variable suffers the usual class imbalance problem (only 15.27% of loans defaulted), which is essential to take into account in the design of the experiments.
Table 1. Variable description.
Variable
Description
Quantitative variables
revenue
Borrower’s self-declared annual income during registration.
dti_n
Indebtedness ratio for obligations excluding mortgage. Monthly information.
loan_amnt
Amount of credit requested by the borrower.
fico_n
Credit bureau score. Defined between 300 and 850, reported by Fair Isaac Corporation as a summary risk measure based on historical credit information reported at the time of application.
Categorical variables
emp_length
Employment length of the borrower categorized into 12 categories, including the no information category.
purpose
Credit purpose category for the loan request.
home_ownership
Homeownership status provided by the borrower.
addr_state
Borrower’s residence state from the USA.
Textual variable
desc
Description of the credit request provided by the borrower.
Regarding the variables that represent the features of the loans, we will consider that we are developing a granting model and, consequently, we will only use those variables known at the time of application. Table 1 shows the descriptions of the variables and we explain them below.
As for the quantitative variables, the information from the Fair Isaac Corporation credit bureau (FICO) in the original dataset is given by a minimum and maximum range of limits to which the borrower’s FICO belongs at loan origination. However, we average these two values to have a single indicator of the creditworthiness of potential borrowers resulting in our fico_n variable. For the case of the debt variable, dti_n is estimated from the original dataset variables as the ratio calculated from the total debts of the co-borrowers over the total debt obligation divided by the combined monthly income of the co-borrowers.
Regarding the categorical variables, we opted to merge the categories ‘other’, ‘none’, and ‘any’ into a unified category labeled ‘other’ for the home_ownership variable. This decision was made due to a lack of clear differentiation among these options, coupled with their similar default percentages and their relatively low percentages of occurrences. The emp_length variable was treated as categorical rather than numerical since it includes categories for ‘no information’ and for ‘more than ten years’.
For the textual variable, we carried out an exhaustive work of text cleaning. First of all, we removed all those descriptions that contained the default description provided by Lending Club on its web form (“Tell your story. What is your loan for?”). Moreover, we removed the prefix “Borrower added on DD/MM/YYYY $>$ from the descriptions, as we did not want any temporal background on them. Finally, as these descriptions came from a web form, we substituted all the HTML elements by their character (e.g. ‘&amp;’ was substituted by ‘&’, ‘&lt;’ was substituted by ‘$<$’, etc.).
Table 2. Exploratory data analysis. Quantitative variables.
Variable Statistic Non-Default Default Total
Mean \$ 73,570.69 \$ $6$ 72,447.84
Median \$ 64,$0$ 58,$0$ 62,000.00
revenue SD \$ 54,944.60 \$ 40,731.98 \$ 53,086.79
KS D-Test 0.09111Significant at the 0.01 level.
Mean 16.15 17.67 16.38
Median 15.88 17.70 16.16
dti_n SD 7.50 7.53 7.53
KS D-Test 0.09111Significant at the 0.01 level.
Mean \$ 13,799.25 \$ 15,111.44 \$ 13,999.66
Median \$ 12,$0$ 14,$0$ 12,000.00
loan_amnt SD \$ 7,931.55 \$ 8,363.19 \$ 8,012.86
KS D-Test 0.08111Significant at the 0.01 level.
Mean 705.22 694.20 703.54
Median 697.00 687.00 697.00
fico_n SD 33.32 27.48 32.74
KS D-Test 0.15111Significant at the 0.01 level.
In Table 2 we show the quantitative variables and use the Kolmogorov-Smirnov test to compare the empirical cumulative probability distributions between the Default and Non-default loans. According to these results, defaulted loans are characterized by lower revenue, higher indebtedness (dti_n), higher requested amount (loan_amnt), and lower FICO (fico_n), being the differences significant at the 0.01 level.
Table 3. Exploratory data analysis. Categorical variables.
Variable Category Count Rel. Freq. Default Rate Chi Test
MORTGAGE 60,796 51.05% 14.14%
OTHER 143 0.12% 20.98%
home_ownership OWN 9,582 8.05% 15.69% 131.08111Significant at the 0.01 level.
RENT 48,580 40.79% 16.60%
$<$ 1 year 9,548 8.02% 14.83%
1 year 7,803 6.55% 14.39%
2 years 10,960 9.20% 14.68%
3 years 9,370 7.87% 14.18%
4 years 7,561 6.35% 14.56%
5 years 9,019 7.57% 14.76%
emp_lenght 6 years 7,271 6.10% 16.04% 104.96111Significant at the 0.01 level.
7 years 6,638 5.57% 15.59%
8 years 5,374 4.51% 15.39%
9 years 4,356 3.66% 15.79%
10+ years 36,287 30.47% 15.41%
NI 4,914 4.13% 19.78%
car 1,884 1.58% 9.61%
credit card 25,051 21.03% 12.81%
debt consolidation 68,372 57.41% 16.14%
educational 265 0.22% 16.98%
home improvement 7,170 6.02% 12.93%
house 805 0.68% 15.78%
purpose major purchase 3,062 2.57% 10.65% 568.47111Significant at the 0.01 level.
medical 970 0.81% 17.32%
moving 768 0.64% 14.84%
other 6,361 5.34% 17.69%
renewable energy 127 0.11% 19.69%
small business 2,518 2.11% 26.41%
vacation 561 0.47% 16.22%
wedding 1,187 1.00% 12.47%
Similarly, in Table 3 we show how the categories are distributed in each categorical variable and the default rate for each category. The addr_state is not included because it has 50 categories (one for each USA state). The table also shows whether there is a significant dependence between the target variable and the categorical variables at the 0.01 significance level. We can see that the dependence is significant for all the variables, including the addr_state not reported in the table (test value of 211.12).
In the home_ownership variable, the ‘OTHER’ category shows the highest risk (20.98%) but a small frequency (0.12%), while the ‘MORTGAGE’ category is the most frequent (51.05%) and the least risky (14.14%) one. In the emp_length variable, the category that denotes no information (‘NI’) has the highest risk (19.78%), but also the lowest frequency (4.13%). In general, the risk in employment length can be divided into two groups with similar default rates, over and below or equal to five years, and curiously the risk of the first group is slightly higher. The categories are not perfectly ordered. These facts support the inclusion of this variable as a categorical variable using one-hot encoding. Finally, the most frequent purpose is ‘debt consolidation’, constituting 57% of the loans, which has a default rate of 16.14%. Notably, the riskiest purpose is ‘small business’, with a 26.41% default rate. Conversely, ‘car’ loans demonstrate the lowest risk, with a mere 9.61% default rate. This striking divergence in default rates across diverse purposes underscores a significant variability in risk within the various loan purposes.
Table 4. Exploratory data analysis. Textual variable.
Variable Statistic Non-Default Default Total
Mean 36.72 35.49 36.54
Median 24.0 22.0 24.0
Word count SD 46.62 48.78 46.96
KS D-Test 0.03111Significant at the 0.01 level.
Mean 66.70 66.31 66.64
Median 73.88 74.19 74.02
Readability SD 32.87 35.55 33.29
KS D-Test 0.02111Significant at the 0.01 level.
Mean 0.0964 0.0909 0.0956
Median 0.0367 0.0 0.0320
Polarity SD 0.1685 0.1699 0.1687
KS D-Test 0.04111Significant at the 0.01 level.
Mean 0.3193 0.3029 0.3168
Median 0.3635 0.3333 0.3589
Subjectivity SD 0.2542 0.2595 0.2551
KS D-Test 0.04111Significant at the 0.01 level.
Regarding the textual description of the loan (desc variable), Table 4 shows some metrics to characterize it. There is a one-word difference in the average word count between the descriptions of defaulted and non-defaulted obligations. The readability was calculated using the Flesch Reading Ease Score [^2] Calculated with Textstat (Python library). Source: https://github.com/textstat/textstat., which indicates the approximate educational level required for comfortable comprehension of a given text (higher scores denote greater ease of reading). The texts in both categories have scores around 66, signifying that they can be readily comprehended by students aged 13 to 15. Additionally, we analyzed the average polarity and average subjectivity [^3] Calculated with TextBlob (Python library). Source: https://github.com/sloria/textblob.. The polarity, ranging from -1 to 1 to denote negative or positive sentiment, was observed to be approximately 0.1 in both cases, suggesting a subtle positive sentiment. On the other hand, subjectivity, measuring the presence of judgments and opinions on a scale from 0 to 1, exhibited values close to 0.31 in both categories. This indicates that while the texts in both cases maintain a generally objective tone, there is a discernible inclusion of some judgments or opinions.
Although the distinctions in these metrics between the default and non-default categories are subtle, their significance is confirmed by the Kolmogorov-Smirnov test. Consequently, it is pertinent to incorporate linguistic aspects into credit risk modeling. Our approach for extracting information from the descriptions relies on leveraging LLMs capable of encompassing not just linguistic nuances but also capturing content details. We elaborate on our methodology below.
5. Description of the methodology
This study seeks to assess the enhancement of default prediction by integrating textual descriptions of loans preprocessed through an LLM. Specifically, we employ transfer learning to enable the LLM to generate a score reflecting the probability of loan default, leveraging information derived from the loan’s textual description.
As our baseline, we employ a machine learning classifier that utilizes the complete set of variables available during the loan application process, encompassing both quantitative and categorical variables. Subsequently, we enhance the classifier by incorporating the LLM score—representing default risk inferred from processing the textual description with the LLM—as an additional input variable and proceed to evaluate the resulting differences. The data and code of our experiments are publicly available [^4] Link will be included in final version.
The key components of the experiment are elaborated upon in this section.
5.1. Tuning the classifier
The classification algorithm used was XGBoost (Chen and Guestrin, 2016), which is trained using a stratified k-fold cross-validation with k being equal to 5. A genetic algorithm (Holland, 1992) was employed to fine-tune the hyperparameters. The fitness value of each individual was the average balanced accuracy (BACC) in the 5 validation sets of the cross-validation. We use BACC as it accounts for the imbalanced nature of the dataset. In preliminary experiments, we also considered the area under the receiver operating characteristic (AUROC), which measures the model’s ability to discriminate between positive and negative examples regardless of the classification threshold chosen. However, we observed that the resulting XGBoost classifiers produced poor BACC values (similar to those from a naïve classifier that predicts the majority class) when using the standard 0.5 threshold to make the prediction. We also observed that XGBoost classifiers with extremely similar AUROC values produced very different results in terms of BACC. Thus, we decided to use the BACC measure as it resulted in classifiers with a more stable behavior.
In genetic optimization, each individual is characterized by its genes, that is, the considered hyperparameters of the XGBoost. Table 5 shows the hyperparameters together with their respective ranges.
Table 5. Hyperparameters of XGBoost considered in the genetic optimization and their respective ranges.
Parameter Min. value Max. value
ETA (learning rate) 0.001 0.5
max_depth 2 12
gamma 0 10
min_child_weight 0 10
alpha 0.5 10
lambda 0.5 10
subsample 0.7 1
colsample_bytree 0.3 1
scale_pos_weight 0.1 10
n_estimators 2 500
The evolutionary strategy chosen for the optimization is the “Mu plus lambda” approach, denoted as $\mu+\lambda$, where $\mu$ represents the number of individuals to select for the next generation, and $\lambda$ indicates the number of children to produce at each generation. Unlike traditional approaches where children often replace parents, the $\mu+\lambda$ strategy involves adding both children and parents to produce the next generation. In the context of this research, we set $\mu$ to 150 and $\lambda$ to 150. This configuration, considering the addition of both parents and offspring, results in generations comprising a total of 300 individuals.
Initially, pairs of parents are chosen through tournament selection with a tournament size of 2. Subsequently, the children are generated employing a two-point crossover technique on the parents’ chromosomes with an 80% probability and applying a random resetting mutation with a 20% probability. The random resetting mutation implies that each gene of every child has a 20% chance of acquiring a new random value within its defined range. To create an offspring of 150 children, this selection, crossover, and mutation process is repeated 75 times. Finally, we combine the 150 children and the 150 parents, resulting in generations of 300 individuals, and select the top 150 ($\mu$) according to their fitness value to pass to the next generation.
The evolutionary process consists of 20 iterations, thereby generating a total of 3,000 individuals—each representing a distinct hyperparameter configuration. From this pool of configurations, the one exhibiting the highest fitness is ultimately chosen as the optimal outcome.
5.2. Generating a default score with BERT
In this section, we delineate the methodology employed to generate a default score based on the textual description of the loan. We initiate the process by applying transfer learning utilizing an LLM, specifically BERT in our case. The fine-tuned BERT model produces an outcome within the range of 0 to 1, offering a nuanced indicator rather than a binary classification. This subtle indicator is subsequently integrated into a classifier along with other input variables to predict the likelihood of loan default. The subsequent steps in this process are detailed next.
5.2.1. Transfer learning to produce the default score
The BERT model used [^5] Source: https://tfhub.dev/tensorflow/bert_en_uncased_L-12_H-768_A-12/4 has the following characteristics: L=12 hidden layers (i.e., Transformer encoder blocks), a hidden size of H=768, and A=12 attention heads (enabling its self-attention mechanism to operate concurrently in 12 distinct patterns). Consequently, the output generated by the BERT model are embeddings of size 768. For the transfer learning process, TensorFlow HUB was used due to its seamless integration with subsequent neural network layers.
To leverage the complex language representation of BERT, we decided to freeze the weights of all BERT hidden layers except the final one. Additionally, we augmented the BERT model by introducing supplementary layers post-embeddings to enrich its adaptability for our specific task and generate the default score. This architectural strategy aims to avoid catastrophic forgetting [^6] A common issue during retraining of pre-trained models that may lead to the deterioration of the original language representation., allowing the resulting model to specifically learn our task without compromising the previously acquired knowledge (Biesialska et al., 2020).
We also explored various configurations for the extra layers to determine the one that produces the best results. Specifically, we explored the 126 configurations that result from combining the following options:
  • Using a first extra dense layer of 128, 256, or 512 neurons.
  • Using or not a second extra dense layer of 128 neurons.
  • Adding a dropout layer before or after all the extra layers.
    • Considering a dropout percentage of 0%, 0.10%, 0.20% or 0.30% for all the dropout layers.
  • Using a learning rate of 0.001, 0.0001, or 0.00001.
All the internal hidden layers are dense layers using the ReLU activation function [^7] The Rectified Linear Unit (ReLU) activation function outputs the maximum of zero and the input value, “activating” the neuron if the input is positive.. Additionally, to obtain the probability of belonging to the default class, the last layer of the neural network was configured with a single neuron and a sigmoid activation function [^8] The sigmoid activation function introduces non-linearity and maps the input values to a range between 0 and 1, facilitating binary classification tasks..
The neural network is trained to predict the loan outcome (default or non-default) based on the textual description as input. As a loss function, we use the weighted binary cross-entropy, which measures the difference between the predicted probabilities and the true binary labels of the input data taking into account class weights [^9] Considering class weights is crucial given the imbalanced nature of our dataset.. The training was set up with a batch size of 64 and trained for 25 epochs, with an early stopper of 3 epochs.
5.2.2. Avoiding data leakage in cross-validation
As previously mentioned, our BERT model generates a default probability, which is then integrated into the classifier as an additional quantitative variable. It is important to note that when computing the BERT default probability, we must replicate the exact folds used in the k-fold cross-validation of the boosting algorithms. In each iteration, boosting algorithms are trained using the BERT_score variable. It is essential to prevent BERT from training with validation data from a specific fold to avoid distorting the model’s true performance. Allowing this would incorporate BERT predictions from its training phase in the fold’s validation data, which doesn’t accurately represent the model’s real-world default prediction ability.
To avoid this data-leakage problem, we use the data as shown in Figure 1. This diagram consists of five steps:
  1. (1)
    Description extraction: Textual descriptions are extracted from the original dataset.
  2. (2)
    Folds generation: The exact same folds as in the boosting algorithms are generated for the textual descriptions. This is done by dividing the data into a train set (green color) and a test set (purple color). Each fold gets a different test set.
  3. (3)
    Optimization of the neural network architecture: The train set obtained in the previous step is divided into a 70% train subset (light green color) and a 30% test subset (dark green color). The neural network is trained with the previously mentioned configurations on the training subset (light green color), and is tested by predicting the test subset (dark green color).
  4. (4)
    Default prediction: The optimal configuration obtained in step 3 (lower value of weighted binary cross-entropy in the test subset) is trained with the train and test subsets (light and dark green colors) and is used to predict the default probabilities of the test set obtained in step 2 (purple color). These predictions of unseen data will be the BERT_score values of the current fold.
  5. (5)
    BERT_score integration: Once the BERT_score values of all the folds are generated in step 4, they are incorporated into the original dataset as a quantitative variable.
Figure 1. BERT training diagram.
5.2.3. Result of the transfer learning process
As explained before, we explored 126 neural network configurations in each of the 5-fold cross-validation. The resulting optimal configuration for each validation fold and its loss score are shown in Table 6. Interestingly, the optimal combination of parameters varies across the different folds, and the only parameter that remains constant is the use of the second dense layer. Despite this parameter heterogeneity, the test loss values are quite stable, except for the case of fold 1, which has a slightly higher loss.
Table 6. BERT optimal configuration and score by fold.
Neurons in
1st dense layer
Use 2nd dense layer
(128 neurons) [^1] Values ‘Yes’ or ‘No’ indicate whether the optimal configuration has or has not that layer.
Dropout Learning rate Loss  [^2] Weighted binary cross-entropy of the test set predictions.
Fold 0 512 Yes 20% 0.0001 0.1774
Fold 1 512 Yes 0% 0.00001 0.1793
Fold 2 128 Yes 20% 0.0001 0.1776
Fold 3 256 Yes 0% 0.001 0.1775
Fold 4 512 Yes 0% 0.001 0.1773
5.2.4. Inclusion of the BERT score in the classification model
The experiment architecture is shown in Figure 2. First, the loan descriptions are processed by the fine-tuned BERT model. By using a sigmoid activation function in its final layer, the resulting output can be interpreted as the probability of belonging to the positive class. This probability, denoted as BERT_score, represents a default score solely based on the content of the loan description. Finally, this BERT score is included as an input variable in the XGBoost classifier along with quantitative and categorical variables to generate a final default prediction.
Figure 2. Diagram of the experiment architecture.
6. Analysis of the BERT score
Table 7. Loan descriptions with highest (top) and lowest (bottom) BERT score.
BERT score Real value
Description
0.8562 0 (Non-Default) getting a divorce need new apt. with new furniture because she getting everthing.
0.8149 1 (Default) need help my bills. to help pay my medications and some bills.
0.8131 0 (Non-Default) i can pay of some bills for my self because i been helping other people out. i could save more for my family and their need.i have a good job that i am bless with. i am from a large family seem like every one thinking i suppose to help them when i need help my self.i alway belive that the lord will.
0.8051 0 (Non-Default) consolidating our debt makes our life easier live in our means with one solid low monthly payment insted of multiple payment that add up more then what ill be paying with this loan and have a little left to leave in my savings for a rainy day to be honest and thank you for your consideration good dy.
0.8045 0 (Non-Default) To consolidate debt. to pay off dept
0.0735 0 (Non-Default) Debt consolidation with a lower APR.
0.1146 0 (Non-Default) In need of funds to pay off some bills as well as minor improvements to house and yard. I have an extremely secure career, and maintaining my credit worthiness is important to me.
0.1262 0 (Non-Default) Hard working individual with a stable job will use loan proceeds to consolidate outstanding credit cards balances. 1) Net monthly income - \$4,432. 2) All expenses (allocated):. Rent - \$1,124. Utilities- 84. Groceries 293. Auto (including fuel) 201 . Cell Phone 52. Cable/Internet 64. Personal care items 82. Entertainment/dining 93. Sales tax 65. 3) Previously answered. 4) No.
0.1360 0 (Non-Default) This loan is to pay off credit ca.
0.1871 1 (Default) I need money for moving expenses and for a buffer for the first month while I transition into working in my new location. I have successfully paid off two previous Lending Club loans in the past couple of years.
In this section, we analyze the BERT score generated. Table 7 shows the loan descriptions with the lowest and highest BERT scores—representing loans identified by BERT as having the highest and lowest default risks, respectively. At first sight, we can see that those with higher BERT scores correspond to poorly descriptive text containing typos (e.g., “pay of”, “everthing”, “alway”, “belive”) or not following grammar rules. Counter-intuitively, only one of these loans eventually defaulted. However, descriptions associated with lower BERT default predictions are more detailed, providing further insight into the purpose of the loan and emphasizing the borrower’s creditworthiness. Some even include numeric values that depict the financial condition of the borrower.
Figure 3. Default and non-default proportions by BERT score range.
In Figure 3 we analyze the proportion of default and non-default loans (Y axis) across the BERT score range (X axis) in bins of 0.01 parts per unit. In the bar, the blue part represents the percentage of Non-Default, while the orange part is the percentage of default observations [^10] The absence of a bar means that there are no loan descriptions with a BERT score within that range.. The figure shows that, in general, as the BERT score increases, the proportion of defaulted loans also increases. It is crucial to note that few observations fall outside the range of $[0.3,0.7]$, rendering the bars beyond this range less reliable. Nevertheless, as a general trend, a higher BERT score value tends to correspond to an elevated probability of default, which underscores the utility of the BERT score as a valuable risk indicator.
Table 8. Classification performance of BERT binarization at 0.5.
Default Non-default
Model BACC Precision Recall F1 Precision Recall F1
BERT 0.5444 0.1714 0.6896 0.2746 0.8771 0.3993 0.5487
Table 8 shows the classification performance obtained by applying a 0.5 threshold to binarize the BERT score, which obtains a balanced accuracy of 54.4%. The BERT score is not very precise in predicting the default class (17.1%) but retrieves 69% of the instances.
These results show that the BERT model can be successfully fine-tuned to learn the task of predicting the final state of the loan using only the information provided by the borrower in the description field of the application form. In Section 7, we will further contextualize this result in light of the results provided by the XGBoost and different sets of variables.
6.1. Relationship between the BERT score and other variables
Table 9. Correlation coefficients of quantitative variables with the BERT score.
Variable Pearson Spearman
revenue -0.0734* -0.1007*
dti_n 0.0663* 0.0627*
loan_amnt -0.0008 -0.0012
fico_n -0.1293* -0.1293*
* p-value less than 0.01.
In this section, we aim to explore potential relationships between the BERT score and other variables. Table 9 presents the correlation coefficients between the BERT score and the quantitative variables within the dataset. The findings reveal weak yet statistically significant relationships with all variables except for the loan amount. Notably, the most pronounced correlations exist with the FICO score and revenue variables. Both exhibit inverse relationships, indicating that individuals with higher FICO scores and revenues tend to have lower BERT scores, and vice versa. The association between the BERT score and the debt variable (dti_n) is direct, albeit slightly weaker than the other two correlations.
Table 10. Kruskal-Wallis test of categorical variables with the BERT score.
Variable H-statistic
emp_length 1581.67*
purpose 1357.94*
home_ownership 243.20*
addr_state 360.44*
* p-value less than 0.01.
To evaluate the relationship with the categorical variables, Table 10 presents the results of the Kruskal-Wallis test. The Kruskal-Wallis test is a non-parametric statistical test that analyzes whether there are statistical differences in the BERT scores across categories within each categorical variable. The results indicate significant differences in BERT scores among all categorical variables, suggesting a certain level of association between the BERT score and these categorical factors. However, it remains challenging to quantify the strength of this relationship or identify the specific categories with the most robust associations.
Table 11. Correlation coefficients of linguistic features with the BERT score.
Variable Pearson Spearman
Word count -0.1961* -0.2650*
Polarity -0.1036* -0.1473*
Subjectivity -0.1753* -0.1866*
Readability 0.1182* 0.1518*
* p-value less than 0.01.
Finally, in Table 11, we investigate the potential relationship between the BERT score and various linguistic features automatically extracted from the text. We observe statistically significant correlations between the BERT score and all considered linguistic features. Notably, there is a relatively strong inverse correlation with both word count and subjectivity, implying that shorter and more objective texts may exhibit higher BERT scores. This observation, coupled with the correlations outlined in Table 9, suggests that BERT demonstrates a stronger association with linguistic aspects than with the numerical variables characterizing the loan application.
7. Classification results
7.1. Effect of including the BERT score
First, we evaluate the impact of incorporating the BERT score on classification performance. Table 12 compares the baseline experiment, featuring an optimized XGBoost with a genetic algorithm using quantitative and categorical variables, against the same approach augmented by the inclusion of the BERT score. A closer examination of the balanced metrics reveals a marginal enhancement in both BACC (0.6154 vs. 0.6187) and AUC (0.6575 vs. 0.6644), being the latter significant according to the DeLong test (Sun and Xu, 2014) at the 0.01 level. It is crucial to note that the Lending Club dataset exclusively consists of approved loans. This fact poses a substantial challenge to significantly enhance the outcomes in a loan granting model such as ours, since the loans included in the dataset were initially considered favorable by the platform.
Table 12. Classifier performance with and without the BERT score.
Metric Quant. + Categ. var. Quant. + Categ. + BERT score
BACC 0.6154 0.6187
AUC 0.6575 0.6644
F1 0.3266 0.3308
Precision 0.2168 0.2249
Recall 0.6614 0.6360
Accuracy 0.5835 0.6066
Table 13. Agreement and disagreement of the classifier with and without the BERT score.
Quant. + Categ. + BERT
Hits Misses
Quant. + Categ. Hits 63907 5004
Misses 8046 42144
Disagreement % 10.96%
The performance metrics in Table 12 indicate that incorporating the BERT score results in improved precision but diminished recall. These values suggest a discernible impact on the classifier’s behavior with the introduction of the new variable. Table 13 examines whether both classifiers yield the same outcome and whether it is accurate (hit) or inaccurate (miss). Notably, the cells in the inverse diagonal exhibit a significant number of instances, accounting for 11% of the total. When assessing the disparity in the accuracy of both classifiers, the inclusion of the BERT score leads to a modest 2% improvement. Therefore, while the incorporation of the BERT score enhances balanced performance, it does so not only by rectifying misses from the other classifier but also by influencing predictions for other instances.
Table 14 shows an additional experiment in which classifiers only use one kind of input variable. While the classifier using the quantitative variables is clearly the best, the classifier that uses just the BERT score obtains slightly better results than the one using the four categorical variables [^11] The difference in the AUC is significant at the 0.01 level according to the DeLong test.. This result is noteworthy given the well-known effectiveness and the meaningful nature of the qualitative variables.
Table 14. Performance considering only a kind of variables.
Metric Quant. Categ. BERT score
BACC 0.6062 0.5486 0.5490
AUC 0.6457 0.5656 0.5714
F1 0.3192 0.2759 0.2601
Precision 0.2138 0.1746 0.1877
Recall 0.6300 0.6563 0.5153
Accuracy 0.5896 0.4738 0.5724
7.2. Feature importance and explainability
To further analyze the impact of the BERT score, in this section, we analyze whether the models’ use of the input variables changes. Figure 4 shows the feature importance given by XGBoost both with and without the incorporation of the BERT score. Notably, the BERT score emerges as the third most influential variable, accounting for 6% importance. Moreover, the order and magnitude of importance for other variables exhibit shifts. Specifically, the borrower’s annual revenue diminishes in importance following the inclusion of the BERT score, while the significance of the ‘credit card’ purpose increases. This observation underscores the significant influence of the BERT score on the classifier.
(a) Before BERT score incorporation.
(b) After BERT score incorporation.
Figure 4. Feature importances of the XGBoost classifiers.
To gain a deeper understanding of the role of variables, we employ SHAP values (Lundberg and Lee, 2017). Figure 5 provides a visualization of the SHAP values for the 10 variables with the most significant impact on the model’s output, comparing models with and without the inclusion of the BERT score.
(a) Before BERT score incorporation.
(b) After BERT score incorporation.
Figure 5. SHAP values of the XGBoost classifiers.
It is noteworthy that Figure 4(a) resembles a comparable figure presented in (Ariza-Garzón et al., 2020), wherein SHAP values were utilized in a lending model on Lending Club [^12] Our dataset is more constrained, only considering loans with textual descriptions, as explained in Section 4.. Examining Figure 4(b), we find that the BERT score ranks as the fourth variable in terms of impact. A closer inspection of the violin plots reveals that the incorporation of BERT does not substantially alter the way variables influence the target class.
Figure 6. Dependence plot between SHAP values and BERT score.
Figure 6 reveals a direct and linear relationship between the BERT score and the default risk. Notably, this relationship is asymmetric around BERT score values of 0.5; scores below 0.4 correspond to SHAP values ranging between -0.4 and -0.6, leading the model to predict non-default. Conversely, this impact range in the positive case is only reached by BERT scores exceeding 0.7, signifying that only exceptionally high BERT scores serve as strong indicators of default.
7.3. Impact across the purpose categories
Now we delve into the influence of incorporating the BERT score on classification outcomes across various categories of the purpose variable. Leveraging the language comprehension capabilities of LLMs, it is conceivable that the BERT model can provide more accurate characterizations of loan purposes than the purpose variable alone. For instance, the ‘other’ category, being inherently ambiguous, could benefit from a more nuanced understanding of the loan description, potentially leading to improved prediction outcomes.
Figure 7. BACC changes by purpose after including the BERT score.
Figure 7 illustrates the changes in balanced accuracy upon incorporating the BERT score for each loan purpose. Notably, the ‘other’ category shows a modest improvement of 2.11%. However, several other categories exhibit more substantial enhancements, including ‘educational’ (9.22%), ‘moving’ (5.84%), ‘medical’ (3.84%), and ‘small business’ (3.40%). These categories share a commonality—the more detailed specification of purpose or a deeper understanding of the borrower’s situation contributes to a more precise delineation of the default risk. For instance, in education loans, factors like the field of study and the educational institution may play a crucial role in determining employability and salaries, thereby enabling a more accurate assessment of the default risk. The same happens in the case of moving to a new place, having to pay medical expenses, or obtaining funds for a business.
The following examples illustrate instances of loans that were initially misclassified as defaults in the absence of the BERT score but were subsequently correctly predicted as non-defaults after its inclusion:
Purpose: Educational; BERT score: 0.3765
“I’m 25 years old and living in New Orleans. I’m asking for a relatively small amount of money to help me take care of my post-Bac tuition for teacher certification and to help me pay off a credit card. I currently work as a private school teacher making very little money with no benefits (about \$29,000 a year). I have to pay about \$1000 in the coming year for my tuition, and I have to get health insurance ASAP, but it’s hard to do so with no financial help from anyone else. My parents can’t help me because my mother is permanently disabled and my father took a huge pay cut this year.”
Purpose: Moving; BERT score: 0.3843
“Although I can afford payments,due to some recent expenses, I am short on cash flow for an unexpected move. I am, however, looking for a more reasonable alternative to banking rates. I have borrowed from Lending Club before and always paid fully and on time with automatic payments.”
Purpose: Small business; BERT score: 0.3292
“The purpose of this loan is to fund advertising costs for a growing internet business venture. I am a successful sales professional earning an average of over \$250K per year over the last 5 years. My credit scores are strong and I have a documented history of paying all my debts (personal or business related) on time.”
Purpose: Medical; BERT score: 0.3921
“This loan will be used to pay off a Care Credit credit card currently at 21.9% I used the card to pay for a prosthetic limb that my health insurer would not cover.”
In each of these examples, the corresponding BERT scores fall below 0.4. As demonstrated in Figure 6, when the score is below this value, the classifier will have a stronger driver to categorize the loan as non-default. All the examples offer more details on the situation and the purpose of the loan. However, given the black-box nature of our BERT-based approach, the specific factors influencing the BERT score cannot be precisely determined. Furthermore, as we have shown, the score is also related to linguistic factors like tone, or language accuracy, and all the texts shown are accurate, readable, and written in a confident and objective tone. However, previous evidence in the literature offered mixed results about the influence of linguistic factors as default predictors (Gao and Lin, 2015; Wang et al., 2016; Dorfleitner et al., 2016; Siering, 2023).
8. Conclusions
In this paper, we have introduced a novel approach leveraging state-of-the-art natural language processing techniques to enhance credit risk models in P2P lending. Specifically, we employed BERT, a widely adopted LLM, to analyze the loan descriptions provided by borrowers. We successfully fine-tuned BERT to distinguish between defaulted and non-defaulted loans and used the model’s output as a risk score. Incorporating the BERT-based risk score improves the classification performance of a conventional granting model in a statistically significant manner. In fact, employing only this score as an input variable yields superior performance compared to considering the categorical variables alone. These results showcase the potential of LLMs to enhance lending models that rely on traditional variables.
Our approach can be easily applied without the need for manual annotation, which is a time-consuming process that also introduces annotator subjectivity. Once the LLM is trained, it instantly generates the risk score for a loan description. As we have illustrated, the score takes into account both linguistic factors and content aspects that relate to the creditworthiness of the borrower’s situation and purpose. However, the true drivers behind a given score or any potential biases remain uncertain. This lack of transparency and explainability is crucial not only for regulatory compliance but also for fostering trust among borrowers and lenders (ROFIEG, 2019).
These facts underscore the need for further exploration and integration of cutting-edge methodologies to enhance the effectiveness of risk-return assessments in the dynamic landscape of P2P lending or even in traditional financial institutions. Several directions deserve to be explored to refine the predictive capability of the model and the understanding of loan applicants’ conditions.
A potential research avenue involves delving into more advanced LLMs. For instance, encoder-only models like RoBERTa (Liu et al., 2019) could improve the model’s capacity to capture more intricate linguistic patterns. Furthermore, the rise of generative AI has spurred interest in exploring decoder models for classification purposes. Although these models were not originally designed for such tasks, recent advancements have shown promising approximations. For example, CARP (Clue And Reasoning Prompting) (Sun et al., 2023) utilizes in-context learning, generating textual responses conditioned on a provided prompt with a few annotated examples (few-shot learning). Remarkably, this approach improved results in classification tasks without the need for transfer learning and fine-tuning and could be applied to risk scoring.
Another promising area of research is to use these language models in a less opaque way, such as through the use of topic models. While many topic models struggle to generate interpretable topics with large and heavily tailed vocabularies, novel approaches that combine traditional topic modeling and word embeddings (Dieng et al., 2020) can address these challenges. Applying such topic modeling approaches to loan descriptions could facilitate the identification of situations with varying risk levels. These identified topics can then be incorporated as additional variables in the classifier model, enhancing overall predictive performance without compromising the demanded transparency.