### v2

# Credit Risk Meets Large Language Models: Building a Risk Indicator from Loan Descriptions in P2P Lending

Abstract
Peer-to-peer (P2P) lending connects borrowers with lenders through online platforms but faces significant information asymmetry, as lenders often lack sufficient data to assess the creditworthiness of borrowers. This paper addresses this challenge by leveraging BERT (Bidirectional Encoder Representations from Transformers), a Large Language Model (LLM) known for its ability to understand contextual nuances in text, to analyze borrowers’ loan descriptions.
We apply transfer learning to make BERT distinguish between default and non-default loans using the loan descriptions of the Lending Club dataset. The resulting BERT-generated risk score demonstrates solid competence in discriminating between default and non-default loans. Furthermore, integrating the BERT-generated risk score with traditional variables enhances classifier performance, which demonstrates the complementary nature of advanced language model outputs in refining credit risk assessment methodologies.
However, the opacity of LLMs and potential biases highlight the need for transparent regulatory frameworks. This study opens new research avenues in P2P lending and AI, emphasizing the importance of trust and transparency in adopting advanced credit risk models.
keywords:
Credit Risk , P2P Lending , BERT , Transfer Learning , Explainable AI , XGBoost
[…]
[ucm] organization=Facultad de Informática, Universidad Complutense de Madrid, addressline=Calle del Prof. José García Santesmases, 9, city=Madrid, postcode=28040, state=Comunidad de Madrid, country=Spain
[…]
[itc] organization=Instituto de Tecnología del Conocimiento, Universidad Complutense de Madrid, addressline=Calle del Prof. José García Santesmases, 9, city=Madrid, postcode=28040, state=Comunidad de Madrid, country=Spain
{graphicalabstract}
{highlights}
Fine-tuned BERT successfully produces risk scores from loan descriptions.
Integrating BERT score with traditional variables boosts granting model performance.
The approach applies without manual annotation, reducing subjectivity and complexity.
Generating risk scores is instant with standard equipment.
Research is needed to enhance transparency and trust in these LLM-based approaches.
1 Introduction
Peer-to-peer (P2P) lending is a growing phenomenon that allows individuals to engage in direct lending and borrowing transactions, bypassing traditional financial institutions. The process is facilitated through online platforms, where prospective borrowers submit loan applications and potential lenders make informed decisions about where to invest their funds.
An inherent challenge in P2P lending is the presence of information asymmetry, wherein borrowers possess more and often superior information compared to lenders. To address this issue, platforms employ strategies to complement the conventional data provided in loan applications [1]. For instance, borrowers are frequently encouraged to provide a voluntary textual description describing the purpose of the loan and their particular situation. Despite the absence of formal verification, such voluntary disclosures have been observed to stimulate increased bidding activity among lenders. However, lenders may lack the expertise to assess the creditworthiness of borrowers effectively and may be influenced by different factors [2].
Traditional credit scoring models do not harness the rich information embedded in the narratives submitted by loan applicants. Several approaches have tried to incorporate such information into credit risk models. The approaches include the extraction of linguistic metrics [3, 4], the use of topic modeling to characterize the underlying themes [5, 6], or a combination of both [7].
However, these methods have limitations in fully capturing the contextual nuances and semantic intricacies of loan descriptions. This study aims to address this gap by leveraging the capabilities of Large Language Models (LLMs), specifically BERT (Bidirectional Encoder Representations from Transformers) [8], which has shown exceptional performance in understanding and processing complex textual data.
BERT’s bidirectional training and ability to capture context at a granular level make it particularly suitable for tasks requiring deep semantic understanding. It has proven successful in fine-tuning for classification tasks [9], in particular domains such as biomedicine [10] and in specialized tasks like spam classification [11]. In this study, we apply transfer learning to train BERT to distinguish between default and non-default loans based on their textual descriptions.
Our work demonstrates BERT’s capability to effectively leverage descriptive data to generate a risk indicator that accurately classifies defaulted loans. Moreover, we show that integrating the BERT-generated risk score into traditional credit granting models can significantly enhance their predictive performance. This highlights BERT’s potential to improve the accuracy and reliability of credit risk models in P2P lending.
Furthermore, our findings highlight the need for greater transparency in adopting advanced models to build trust among users and regulatory entities. This study paves the way for future research, emphasizing the transformative potential of LLMs like BERT in the field of credit risk assessment.
This paper is organized as follows: Section 2 provides a comprehensive review of related work in credit risk assessment and natural language processing. Section 3 presents an overview of LLMs and the BERT model. Section 4 describes the dataset used, detailing the data preprocessing steps and conducting an in-depth exploratory data analysis. Section 5 outlines the methodology, model architecture, and training procedures employed in integrating BERT into the credit risk assessment framework. Section 6 analyzes the risk score generated by the BERT description processing. Section 7 discusses the results of our experiments, highlighting the improvements achieved by incorporating BERT-based textual analysis. Finally, Section 8 concludes the paper by summarizing key findings, discussing implications, and suggesting avenues for future research in the intersection of NLP and credit risk assessment.
2 Related work
In their comprehensive analysis of risk-return modeling within the P2P lending market [12], the authors identify a discernible trend towards including new sources and types of information to improve risk and profit management models in the P2P market. The sources are very diverse and include transactional data [13], the topology of the lending-borrowing network [14], data from social networks [15], or, more recently, facial features [16]. Among them, the authors identify as a predominant trend the inclusion of textual data taken from statements describing the purpose of the loan.
In a pioneer work exploring the impact of textual factors on peer-to-peer lending [17], the authors analyze P2P loans including manually annotated narrative aspects, such as trustworthiness, economic hardship, hard work, success, morality, and religiosity. These aspects were combined with demographic variables and loan characteristics. Their results highlight that narratives regarding trustworthiness strongly influence decision-makers, particularly credit lenders, in their loan approval process. Additionally, some of these narratives play a substantial role in subsequent loan performance.
However, most subsequent studies typically use text mining or artificial intelligence methods to extract linguistic features or loan description topics. Regarding the use of linguistic features, the authors in [3] use machine learning and text mining techniques to quantify and extract linguistic features (e.g., readability, positivity, objectivity, and deception cues), and then build both explanatory econometric models and predictive models using such features. They find that they can indeed reflect borrowers’ creditworthiness and predict loan default. They also use a panel of investors and confirm that investors indeed value texts written by borrowers, but that they can also be deceived by some of the deception cues well established in the literature. Similarly, in [4], the authors include linguistic factors and the presence of social and emotional keywords and evaluate their impact on two European platforms. They found that text-derived variables influence the probability of funding, but not the probability of default. In [18], linguistic statistical features and abstract text features (including deception, subjectivity, sentiment, readability, personality, and mindset) are used to characterize text descriptions. They compare the performance of different classifiers based on the textual features and conclude that their performance is close to that of the classifiers using traditional financial features, but that adding textual features can improve the performance of the whole credit risk evaluation system.
As for topic modeling, the Latent Dirichlet Allocation model (LDA) has been widely used. In [19], the authors use LDA to extract six topics from the loan descriptions whose meanings are obvious: assets, income and expenses, work, family, business, and agriculture. They also consider the number of characters in the descriptive text. They conclude that soft (qualitative) information can improve the performance of loan default prediction compared to existing methods based only on hard (quantitative) information and that soft features have a significant ability to discriminate loan defaults. Similarly, in [20], an LDA topic model is used to classify the loan titles into six purposes. Their findings reveal that the stated purpose significantly influences a borrower’s chances of securing financing. Notably, ambiguous titles—where borrowers fail to clearly articulate the loan’s purpose—substantially diminish the likelihood of loan approval. In [5], Xia et al. used a keyword clustering algorithm for automatic topic extraction. Their method combines keyword extraction based on term frequency-inverse document frequency (TF-IDF) with word embeddings generated by the Word2Vec neural network model [21]. Analysis of three real-world datasets demonstrated that incorporating these topic variables significantly enhanced predictive accuracy compared to relying solely on traditional information.
More recently, Siering [7] investigated the effect of both aspects: topics and linguistic features. To extract topics, the author utilized a financial text analysis procedure [22] to build a domain dictionary. The identified topics describe the purpose of the loan, as well as the borrowers’ request for help, reliability, or appreciation. These topics were operationalized as binary indicator variables. The author employs text mining to create variables that measure aspects such as polarity, active orientation, readability, average sentence length, and word count. These variables are then used as input in a logistic regression. The results indicate that both linguistic and content-based factors contribute to predicting loan default probability, with the latter demonstrating greater significance. Analysis of variable contributions shows that certain factors increase or decrease default likelihood. Notably, expressions of reliability correlate positively with loan repayment.
Yet, a significant gap becomes apparent in applying state-of-the-art natural language processing techniques, including deep learning methods. In their work [6], Zhang et al. studied the transformer encoder’s ability to capture textual features from loan descriptions. These characteristics, alongside the traditional hard features derived from loan applications, were inputted into a neural network to predict the likelihood of loan default. The study highlights the effectiveness of transformers, showing better results when including textual loan descriptions compared to models that do not consider them. In this work, we explore using encoder-based LLMs, such as BERT, which has successfully improved classification in other fields.
3 An overview of Large Language Models and BERT
Large Language Models (LLMs) are built upon the Transformer architecture [23], leveraging attention mechanisms to enhance language comprehension. LLMs can be broadly categorized into three primary families, each distinguished by its architecture:
  • 1.
    Encoder-only, widely employed for language comprehension tasks such as text classification, named entity recognition, and extractive question answering. The most famous example is BERT [8], which will be explained in more detail below.
  • 2.
    Decoder-only, designed for generative tasks, exemplified by the well-known GPT models [24, 25]. It is employed in various tasks, including question answering [26], text summarization [27], and programming code generation [28].
  • 3.
    Encoder-decoder models, suited for tasks demanding both language understanding and generation, such as language translation or text summarization. The most influential models are BART [29] and T5 [30].
The selection of the appropriate architecture hinges on the specific requirements of the intended task. Whether it be the nuanced comprehension of language, creative text generation, or the synthesis of both, the versatility of LLMs offers a tailored solution for diverse applications.
We will focus on BERT (Bidirectional Encoder Representations from Transformer), which is a Transformer-based language model introduced by Google researchers in 2018 [8]. BERT’s architecture consists of a stack of encoders from the Transformer model. The bidirectional nature of BERT is key, as it considers both the left and right context of each word, enhancing its ability to understand context-dependent meanings and to be effective in language understanding tasks. Numerous studies have consistently shown that BERT is the most effective linguistic model for various of these tasks [31, 32]. Notably, BERT has 340 million parameters, while the widely recognized GPT-3 model has 175 billion, making BERT 514 times smaller than GPT-3 [25]. Given this significant size difference, BERT can be operated on standard home equipment for model inference, which greatly simplifies its use in practical scenarios. In contrast, GPT, built with a Transformer decoder stack, not only demands much more powerful equipment but is suited for language generation tasks.
BERT stands as a milestone whose success has spurred the development of a diverse family of models that build upon its architecture. Some versions aim to achieve similar performance while having a smaller number of parameters, such as DistilBERT (a distilled version of BERT) [33], or ALBERT (A Little BERT) [34]. Others are adaptations to other languages such as CamemBERT [35] to French or BETO to Spanish [36]. Other proposals aimed to improve upon BERT by modifying some design decisions when pretraining BERT and also training the model longer, as in the case of RoBERTa (Robustly optimized BERT approach) [37], which resulted in improved contextualized representations and enhanced language understanding.
To further elucidate the role of BERT in specialized applications, it is crucial to understand its capacity for transfer learning and fine-tuning. Transfer learning involves using a pre-trained model like BERT, which has initially learned general language patterns from a large corpus to a specific task or dataset. This technique allows us to take advantage of the rich linguistic representations without needing extensive computation from scratch. Fine-tuning involves adjusting the pre-trained model’s parameters to capture the nuances of the target task or application field by further training with new instances from the new context. For example, BioBERT is a BERT model fine-tuned for biomedical text mining tasks like named entity recognition and question answering [10]. Other adaptations have targeted text classification and sentiment analysis in specific datasets [9, 38].
Our research aims to develop a credit risk model by leveraging a BERT-based model. Specifically, we will fine-tune BERT for producing a predictive score indicative of the likelihood of loan default using the textual descriptions of loans provided by borrowers in their application forms. We will assess the scoring effectiveness in a granting model using the well-established dataset from the Lending Club P2P lending platform.
4 Description of the dataset
We have used a public data set of the P2P lending company Lending Club [^1] https://www.kaggle.com/wordsforthewise/lending-club, which is widely used in credit risk publications and the most widely used when dealing with the P2P market [39, 12]. However, instead of using the original dataset, which includes 2,260,699 loans granted by the company between 2007 and 2018, we will use a version modified for proposing granting models [40], used in [41, 42]. Since granting models determine which loans will be fully repaid, its estimation needs loans whose final status is known, that is, that were either fully repaid or defaulted. Thus, the dataset excludes loans in transitory states (in a grace period, late, etc.) and loans with no information on income and indebtedness which is essential to compute the input variables, resulting in 1,347,681 instances.
Additionally, the original dataset contains variables detailing the loan’s lifecycle and other post-application aspects (e.g., the interest rate). In contrast, our version only includes variables available at the time of application, which are those utilized by granting models.
Loan descriptions were inconsistently available, appearing only for certain loans between April 2008 and March 2014. To accurately assess the impact of textual descriptions on default prediction, our analysis focuses solely on the 119,101 loans that include the desc variable. Kolmogorov-Smirnov and chi-square tests were applied to quantitative and categorical variables to assess potential bias from filtering. The lack of significant differences indicates that the filtered dataset is representative of the original dataset.
Table 1: Variable description.
Variable Description
Quantitative variables
revenue Borrower’s self-declared annual income during registration.
dti_n Indebtedness ratio for obligations excluding mortgage. Monthly information.
loan_amnt Amount of credit requested by the borrower.
fico_n Credit bureau score. Defined between 300 and 850, reported by Fair Isaac Corporation as a summary risk measure based on historical credit information reported at the time of application.
Categorical variables
emp_length Employment length of the borrower categorized into 12 categories, including the no information category.
purpose Credit purpose category for the loan request.
home_ownership Homeownership status provided by the borrower.
addr_state Borrower’s residence state from the USA.
Textual variable
desc Description of the credit request provided by the borrower.
In the dataset, the target variable suffers the usual class imbalance problem (only 15.27% of default), which will be considered in the design of the experiments. Table 1 shows the input variables of our granting model, which are explained below.
As for the quantitative variables, the Fair Isaac Corporation credit bureau (FICO) information in the original dataset is given by a minimum and maximum range of limits to which the borrower’s FICO belongs at loan origination. However, we average these two values to have a single indicator of the creditworthiness of potential borrowers resulting in our fico_n variable. For the case of the debt variable, dti_n is estimated from the original dataset variables as the ratio calculated from the total debts of the co-borrowers over the total debt obligation divided by the combined monthly income of the co-borrowers.
Regarding the categorical variables, we merged the categories ‘other’, ‘none’, and ‘any’ into a unified category labeled ‘other’ for the home_ownership variable. This decision was made due to a lack of clear differentiation among these options, coupled with their similar default percentages and their relatively low percentages of occurrences. The emp_length variable was treated as categorical rather than numerical since it includes categories for ‘no information’ and for ‘more than ten years’.
For the textual variable, we carried out an exhaustive work of text cleaning. First, we removed all those descriptions that contained the default description provided by Lending Club on its web form (“Tell your story. What is your loan for?”). Moreover, we removed the prefix “Borrower added on DD/MM/YYYY $>$ from the descriptions, as we did not want any temporal background on them. Finally, as these descriptions came from a web form, we replaced all HTML entities with their corresponding characters (e.g. ‘&amp;’ was substituted by ‘&’, ‘&lt;’ was substituted by ‘$<$’, etc.).
Table 2: Exploratory data analysis. Quantitative variables.
Variable Statistic Non-Default Default Total
revenue Mean \$ 73,570.69 \$ $6$ 72,447.84
Median \$ 64,$0$ 58,$0$ 62,000.00
SD \$ 54,944.60 \$ 40,731.98 \$ 53,086.79
KS D-Test 0.09*
dti_n Mean 16.15 17.67 16.38
Median 15.88 17.70 16.16
SD 7.50 7.53 7.53
KS D-Test 0.09*
loan_amnt Mean \$ 13,799.25 \$ 15,111.44 \$ 13,999.66
Median \$ 12,$0$ 14,$0$ 12,000.00
SD \$ 7,931.55 \$ 8,363.19 \$ 8,012.86
KS D-Test 0.08*
fico_n Mean 705.22 694.20 703.54
Median 697.00 687.00 697.00
SD 33.32 27.48 32.74
KS D-Test 0.15*
* p-value less than 0.01.
Table 2 presents the quantitative variables and the results of the Kolmogorov-Smirnov test, which was used to compare the empirical cumulative distribution functions of Default and Non-default loans. According to these results, defaulted loans are characterized by lower revenue, higher debt-to-income ratio (dti_n), higher requested amount (loan_amnt), and lower FICO scores (fico_n), being the differences significant at the 0.01 level.
Table 3: Exploratory data analysis. Categorical variables.
Variable Category Count Rel. Freq. Default Rate Chi Test
home_ownership MORTGAGE 60,796 51.05% 14.14% 131.08*
OTHER 143 0.12% 20.98%
OWN 9,582 8.05% 15.69%
RENT 48,580 40.79% 16.60%
emp_lenght $<$ 1 year 9,548 8.02% 14.83% 104.96*
1 year 7,803 6.55% 14.39%
2 years 10,960 9.20% 14.68%
3 years 9,370 7.87% 14.18%
4 years 7,561 6.35% 14.56%
5 years 9,019 7.57% 14.76%
6 years 7,271 6.10% 16.04%
7 years 6,638 5.57% 15.59%
8 years 5,374 4.51% 15.39%
9 years 4,356 3.66% 15.79%
10+ years 36,287 30.47% 15.41%
NI 4,914 4.13% 19.78%
purpose car 1,884 1.58% 9.61% 568.47*
credit card 25,051 21.03% 12.81%
debt consolidation 68,372 57.41% 16.14%
educational 265 0.22% 16.98%
home improvement 7,170 6.02% 12.93%
house 805 0.68% 15.78%
major purchase 3,062 2.57% 10.65%
medical 970 0.81% 17.32%
moving 768 0.64% 14.84%
other 6,361 5.34% 17.69%
renewable energy 127 0.11% 19.69%
small business 2,518 2.11% 26.41%
vacation 561 0.47% 16.22%
wedding 1,187 1.00% 12.47%
* p-value less than 0.01.
Similarly, Table 3 displays the distribution of categories within each categorical variable and the corresponding default rates. The addr_state variable is excluded due to its 50 categories, one for each U.S. state. The table also indicates whether there is a significant dependence between the target variable and the categorical variables at the 0.01 significance level. The results show significant dependence for all variables, including the addr_state variable not reported in the table (test value of 211.12).
In the home_ownership variable, the ‘OTHER’ category shows the highest risk (20.98%) but a small frequency (0.12%), while the ‘MORTGAGE’ category is the most frequent (51.05%) and the least risky (14.14%) one. In the emp_length variable, the category that denotes no information (‘NI’) has the highest risk (19.78%), but also the lowest frequency (4.13%). In general, employment length can be categorized into two groups with comparable default rates: those with employment lengths of five years or less and those with more than five years. Interestingly, the risk is slightly higher in the group with more than five years of employment. The categories are not perfectly ordered, which supports the use of one-hot encoding to treat this variable as categorical. Finally, the most frequent purpose is ‘debt consolidation’, constituting 57% of the loans, which has a default rate of 16.14%. Notably, the riskiest purpose is ‘small business’, with a 26.41% default rate. Conversely, ‘car’ loans demonstrate the lowest risk, with a mere 9.61% default rate. This striking divergence in default rates across diverse purposes underscores a significant variability in risk within the various loan purposes.
Table 4: Exploratory data analysis. Textual variable.
Variable Statistic Non-Default Default Total
Word count Mean 36.72 35.49 36.54
Median 24.0 22.0 24.0
SD 46.62 48.78 46.96
KS D-Test 0.03*
Readability Mean 66.70 66.31 66.64
Median 73.88 74.19 74.02
SD 32.87 35.55 33.29
KS D-Test 0.02*
Polarity Mean 0.0964 0.0909 0.0956
Median 0.0367 0.0 0.0320
SD 0.1685 0.1699 0.1687
KS D-Test 0.04*
Subjectivity Mean 0.3193 0.3029 0.3168
Median 0.3635 0.3333 0.3589
SD 0.2542 0.2595 0.2551
KS D-Test 0.04*
* p-value less than 0.01.
Regarding the textual description of the loan (desc variable), Table 4 shows some metrics to characterize it. There is a one-word difference in the average word count between the descriptions of defaulted and non-defaulted obligations. The readability was calculated using the Flesch Reading Ease Score [^2] Calculated with Textstat (Python library). Source: https://github.com/textstat/textstat., which indicates the approximate educational level required for comfortable comprehension of a given text (higher scores denote greater ease of reading). The texts in both categories have scores around 66, signifying that they can be readily comprehended by students aged 13 to 15. Additionally, we analyzed the average polarity and average subjectivity [^3] Calculated with TextBlob (Python library). Source: https://github.com/sloria/textblob.. The polarity, ranging from -1 to 1 to denote negative or positive sentiment, was observed to be approximately 0.1 in both cases, suggesting a subtle positive sentiment. On the other hand, subjectivity, measuring the presence of judgments and opinions on a scale from 0 to 1, exhibited values close to 0.31 in both categories. This indicates that while the texts in both cases maintain a generally objective tone, there is a discernible inclusion of some judgments or opinions.
Although the distinctions in these metrics between the default and non-default categories are subtle, their significance is confirmed by the Kolmogorov-Smirnov test. Consequently, it is pertinent to incorporate linguistic aspects into credit risk modeling. Our approach for extracting information from the descriptions relies on leveraging LLMs capable of encompassing not just linguistic nuances but also capturing content details. We elaborate on our methodology below.
5 Description of the methodology
This study assesses the enhancement of default prediction by integrating textual descriptions of loans preprocessed through an LLM. Specifically, we employ transfer learning and fine-tuning to enable the LLM to generate a score reflecting the probability of loan default, leveraging information derived from the loan’s textual description.
As our baseline, we employ a machine learning classifier that utilizes the complete set of variables available during the loan application process, encompassing both quantitative and categorical variables. Subsequently, we enhance the classifier by incorporating the LLM score—representing default risk inferred from processing the textual description with the LLM—as an additional input variable and proceed to evaluate the resulting differences. The data and code of our experiments are publicly available [^4] Link of the code will be included in final version..
The key components of the experiment are elaborated upon in this section.
5.1 Tuning the classifier
The classification algorithm used was XGBoost [43], which is trained using a stratified k-fold cross-validation with k being equal to 5 and where the instances are shuffled to avoid potential ordering biases in the dataset. A genetic algorithm [44] was employed to fine-tune the hyperparameters. The fitness value of each individual was the average balanced accuracy (BACC) in the 5 validation sets of the cross-validation. We use BACC as it accounts for the imbalanced nature of the dataset. In preliminary experiments, we also considered the area under the receiver operating characteristic (AUROC), which measures the model’s ability to discriminate between positive and negative examples regardless of the classification threshold chosen. However, we observed that the resulting XGBoost classifiers produced poor BACC values (similar to those from a naïve classifier that predicts the majority class) when using the standard 0.5 threshold to make the prediction. We also observed that XGBoost classifiers with extremely similar AUROC values produced very different results in terms of BACC. Thus, we decided to use the BACC measure as it resulted in classifiers with a more stable behavior.
In genetic optimization, each individual is characterized by its genes, that is, the considered hyperparameters of the XGBoost. Table 5 shows the hyperparameters together with their respective ranges.
Table 5: Hyperparameters of XGBoost considered in the genetic optimization and their respective ranges.
Parameter Min. value Max. value
ETA (learning rate) 0.001 0.5
max_depth 2 12
gamma 0 10
min_child_weight 0 10
alpha 0.5 10
lambda 0.5 10
subsample 0.7 1
colsample_bytree 0.3 1
scale_pos_weight 0.1 10
n_estimators 2 500
The evolutionary strategy chosen for the optimization is the “Mu plus lambda” approach, denoted as $\mu+\lambda$, where $\mu$ represents the number of individuals to select for the next generation, and $\lambda$ indicates the number of children to produce at each generation. Unlike traditional approaches where children often replace parents, the $\mu+\lambda$ strategy involves adding both children and parents to produce the next generation. In the context of this research, we set $\mu$ to 150 and $\lambda$ to 150. This configuration, considering the addition of both parents and offspring, results in generations comprising a total of 300 individuals.
Initially, pairs of parents are chosen through tournament selection with a tournament size of 2. Subsequently, the children are generated employing a two-point crossover technique on the parents’ chromosomes with an 80% probability and applying a random resetting mutation with a 20% probability. The random resetting mutation implies that each gene of every child has a 20% chance of acquiring a new random value within its defined range. To create an offspring of 150 children, this selection, crossover, and mutation process is repeated 75 times. Finally, we combine the 150 children and the 150 parents, resulting in generations of 300 individuals, and select the top 150 ($\mu$) according to their fitness value to pass to the next generation.
The evolutionary process consists of 20 iterations, thereby generating a total of 3,000 individuals—each representing a distinct hyperparameter configuration. From this pool of configurations, the one exhibiting the highest fitness is ultimately chosen as the optimal outcome.
5.2 Generating a default score with BERT
In this section, we delineate the methodology employed to generate a default score based on the textual description of the loan. We initiate the process by applying transfer learning utilizing an LLM, specifically BERT in our case. The fine-tuned BERT model produces an outcome within the range of 0 to 1, offering a nuanced indicator rather than a binary classification. This subtle indicator is subsequently integrated into a classifier along with other input variables to predict the likelihood of loan default. The subsequent steps in this process are detailed next.
5.2.1 Transfer learning to produce the default score
The BERT model we utilize [^5] Source: https://tfhub.dev/tensorflow/bert_en_uncased_L-12_H-768_A-12/4 is configured with L=12 hidden layers (i.e., Transformer encoder blocks), each with a size of H=768, and it employs A=12 attention heads. These attention heads enable the model’s self-attention mechanism to process inputs in 12 distinct patterns simultaneously. The output from BERT are embeddings of size 768. For incorporating this model, TensorFlow HUB was selected for its efficient integration with additional neural network layers.
As outlined in Section 3, during the transfer learning phase, we aim to exploit BERT’s advanced language understanding while minimizing the need to learn from scratch. To achieve this, we freeze the weights of all but the last hidden layer of BERT. This approach preserves the model’s pre-trained capabilities and prevents overfitting on our specific dataset, a phenomenon known as catastrophic forgetting [^6] A degradation issue during the retraining of pre-trained models, leading to loss of the original capabilities [45]..
Subsequently, in the fine-tuning stage, only the last BERT layer and the newly added layers are adjusted to better serve our specific task of generating a default score. These layers are optimized to enhance the model’s adaptability to our particular requirements, allowing slight parameter adjustments for improved task-specific performance, while maintaining the general language understanding gained from BERT’s initial pre-training. This strategy effectively balances specialized learning with the retention of valuable pre-trained knowledge.
We also explored various configurations for the extra layers to determine the one that produces the best results. Specifically, we explored the 126 configurations that result from combining the following options:
  • 1.
    Using a first extra dense layer of 128, 256, or 512 neurons.
  • 2.
    Using or not a second extra dense layer of 128 neurons.
  • 3.
    Adding a dropout layer before or after all the extra layers.
    • (a)
      Considering a dropout percentage of 0%, 0.10%, 0.20% or 0.30% for all the dropout layers.
  • 4.
    Using a learning rate of 0.001, 0.0001, or 0.00001.
All the internal hidden layers are dense layers using the ReLU activation function [^7] The Rectified Linear Unit (ReLU) activation function outputs the maximum of zero and the input value, “activating” the neuron if the input is positive.. Additionally, to obtain the probability of belonging to the default class, the last layer of the neural network was configured with a single neuron and a sigmoid activation function [^8] The sigmoid activation function introduces non-linearity and maps the input values to a range between 0 and 1, facilitating binary classification tasks..
The neural network is trained to predict the loan outcome (default or non-default) based on the textual description as input. As a loss function, we use the weighted binary cross-entropy, which measures the difference between the predicted probabilities and the true binary labels of the input data taking into account class weights [^9] Considering class weights is crucial given the imbalanced nature of our dataset.. The training was set up with a batch size of 64 and trained for 25 epochs, with an early stopper of 3 epochs.
5.2.2 Avoiding data leakage in cross-validation
As previously mentioned, our BERT model generates a default probability, which is then integrated into the classifier as an additional quantitative variable. It is important to note that when computing the BERT default probability, we must replicate the exact folds used in the k-fold cross-validation of the boosting algorithms. In each iteration, boosting algorithms are trained using the BERT_score variable. It is essential to prevent BERT from training with validation data from a specific fold to avoid distorting the model’s true performance. Allowing this would incorporate BERT predictions from its training phase in the fold’s validation data, which doesn’t accurately represent the model’s real-world default prediction ability.
To avoid this data-leakage problem, we use the data as shown in Figure 1. This diagram consists of five steps:
  1. 1.
    Description extraction: Textual descriptions are extracted from the original dataset.
  2. 2.
    Folds generation: The exact same folds as in the boosting algorithms are generated for the textual descriptions. This is done by dividing the data into a train set (green color) and a test set (purple color). Each fold gets a different test set.
  3. 3.
    Optimization of the neural network architecture: The train set obtained in the previous step is divided into a 70% train subset (light green color) and a 30% test subset (dark green color). The neural network is trained with the previously mentioned configurations on the training subset (light green color), and is tested by predicting the test subset (dark green color).
  4. 4.
    Default prediction: The optimal configuration obtained in step 3 (lower value of weighted binary cross-entropy in the test subset) is trained with the train and test subsets (light and dark green colors) and is used to predict the default probabilities of the test set obtained in step 2 (purple color). These predictions of unseen data will be the BERT_score values of the current fold.
  5. 5.
    BERT_score integration: Once the BERT_score values of all the folds are generated in step 4, they are incorporated into the original dataset as a quantitative variable.
Figure 1: BERT training diagram.
5.2.3 Result of the transfer learning process
As explained before, we explored 126 neural network configurations in each of the 5-fold cross-validation. The resulting optimal configuration for each validation fold and its loss score are shown in Table 6. Interestingly, the optimal combination of parameters varies across the different folds, and the only parameter that remains constant is the use of the second dense layer. Despite this parameter heterogeneity, the test loss values are quite stable, except for the case of fold 1, which has a slightly higher loss.
Table 6: BERT optimal configuration and score by fold.
Neurons in 1st dense layer Use 2nd dense layer (128 neurons)* Dropout Learning rate Loss** Fold 0 512 True 20% 0.0001 0.1774 Fold 1 512 True 0% 0.00001 0.1793 Fold 2 128 True 20% 0.0001 0.1776 Fold 3 256 True 0% 0.001 0.1775 Fold 4 512 True 0% 0.001 0.1773 * Values ‘True’ or ‘False’ indicate whether the optimal configuration has or has not that layer. ** Weighted binary cross-entropy of the test set predictions.
5.2.4 Inclusion of the BERT score in the classification model
The experiment architecture is shown in Figure 2. First, the loan descriptions are processed by the fine-tuned BERT model. By using a sigmoid activation function in its final layer, the resulting output can be interpreted as the probability of belonging to the positive class. This probability, denoted as BERT_score, represents a default score solely based on the content of the loan description. Finally, this BERT score is included as an input variable in the XGBoost classifier along with quantitative and categorical variables to generate a final default prediction.
Figure 2: Diagram of the experiment architecture.
The fine-tuning of the BERT model was performed on a system equipped with an Intel Core i9-12900KS processor, an NVIDIA GeForce RTX 4090 graphics card (24GB), and 128 GB of RAM. In contrast, the inference process with the fine-tuned model, which involves generating the BERT score for a given loan description, can be conducted on a standard PC. This capability enhances the practicality of deploying our approach in real-world applications.
6 Analysis of the BERT score
6.1 Assessment of the BERT score as a credit risk score
Table 7: Loan descriptions with highest (top) and lowest (bottom) BERT score.
BERT score Real value Description
0.8562 0 (Non-Default) getting a divorce need new apt. with new furniture because she getting everthing.
0.8149 1 (Default) need help my bills. to help pay my medications and some bills.
0.8131 0 (Non-Default) i can pay of some bills for my self because i been helping other people out. i could save more for my family and their need.i have a good job that i am bless with. i am from a large family seem like every one thinking i suppose to help them when i need help my self.i alway belive that the lord will.
0.8051 0 (Non-Default) consolidating our debt makes our life easier live in our means with one solid low monthly payment insted of multiple payment that add up more then what ill be paying with this loan and have a little left to leave in my savings for a rainy day to be honest and thank you for your consideration good dy.
0.8045 0 (Non-Default) To consolidate debt. to pay off dept
0.0735 0 (Non-Default) Debt consolidation with a lower APR.
0.1146 0 (Non-Default) In need of funds to pay off some bills as well as minor improvements to house and yard. I have an extremely secure career, and maintaining my credit worthiness is important to me.
0.1262 0 (Non-Default) Hard working individual with a stable job will use loan proceeds to consolidate outstanding credit cards balances. 1) Net monthly income - \$4,432. 2) All expenses (allocated):. Rent - \$1,124. Utilities- 84. Groceries 293. Auto (including fuel) 201 . Cell Phone 52. Cable/Internet 64. Personal care items 82. Entertainment/dining 93. Sales tax 65. 3) Previously answered. 4) No.
0.1360 0 (Non-Default) This loan is to pay off credit ca.
0.1871 1 (Default) I need money for moving expenses and for a buffer for the first month while I transition into working in my new location. I have successfully paid off two previous Lending Club loans in the past couple of years.
In this section, we evaluate the effectiveness of the BERT score in utilizing loan descriptions to predict default risk. Table 7 presents loan descriptions with the highest and lowest BERT scores, which indicate loans assessed by BERT as having the highest and lowest default risks, respectively. It appears that descriptions with higher BERT scores are often less informative, containing errors such as typos (e.g., “pay of”, “everthing”, “alway”, “belive”) or grammatical issues. Interestingly, only one of these loans ultimately defaulted. Conversely, loan descriptions with lower BERT scores tend to be more detailed and insightful, often including information about the loan’s purpose and the borrower’s creditworthiness, and sometimes even providing numeric data related to the borrower’s financial status.
Figure 3: Default and non-default proportions by BERT score range.
Figure 3 illustrates the distribution of default and non-default loans (Y axis) across the BERT score range (X axis) in bins of 0.01. In the bar chart, the blue segment represents the percentage of non-default loans, while the orange segment denotes the percentage of default loans [^10] Absence of a bar indicates that no loan descriptions fall within that BERT score range.. The figure reveals a general trend where higher BERT scores are associated with a higher proportion of defaulted loans. It is important to note that observations outside the BERT score range of $[0.3,0.7]$ are sparse, which makes the bars in these regions less reliable. Overall, the trend suggests that higher BERT scores are indicative of a greater likelihood of default, highlighting the BERT score’s usefulness as a risk assessment tool.
Table 8: Classification performance of BERT binarization at 0.5.
Default Non-default
Model BACC Precision Recall F1 Precision Recall F1
BERT 0.5444 0.1714 0.6896 0.2746 0.8771 0.3993 0.5487
Table 8 shows the classification performance obtained by applying a 0.5 threshold to binarize the BERT score, which obtains a balanced accuracy of 54.4%. The BERT score is not very precise in predicting the default class (17.1%) but retrieves 69% of the instances.
These results demonstrate that the BERT model can be effectively fine-tuned to predict the final state of the loan using only the information provided by the borrower in the description field of the application form. In Section 7, we will contextualize these findings by comparing them with the results obtained from XGBoost and various sets of variables.
6.2 Assessment of the relationship between the BERT score and other variables
In this section, we aim to explore potential relationships between the BERT score and other variables. Table 9 presents the correlation coefficients between the BERT score and the quantitative variables within the dataset. The findings reveal weak yet statistically significant relationships with all variables except for the loan amount. Notably, the most pronounced correlations exist with the FICO score and revenue variables. Both exhibit inverse relationships, indicating that individuals with higher FICO scores and revenues tend to have lower BERT scores, and vice versa. The association between the BERT score and the debt variable (dti_n) is direct, albeit slightly weaker than the other two correlations.
Table 9: Correlation coefficients of quantitative variables with the BERT score.
Variable Pearson Spearman
revenue -0.0734* -0.1007*
dti_n 0.0663* 0.0627*
loan_amnt -0.0008 -0.0012
fico_n -0.1293* -0.1293*
* p-value less than 0.01.
Table 10: Kruskal-Wallis test of categorical variables with the BERT score.
Variable H-statistic
emp_length 1581.67*
purpose 1357.94*
home_ownership 243.20*
addr_state 360.44*
* p-value less than 0.01.
To evaluate the relationship with the categorical variables, Table 10 presents the results of the Kruskal-Wallis test. The Kruskal-Wallis test is a non-parametric statistical test that analyzes whether there are statistical differences in the BERT scores across categories within each categorical variable. The results indicate significant differences in BERT scores among all categorical variables, suggesting a certain level of association between the BERT score and these categorical factors. However, it remains challenging to quantify the strength of this relationship or identify the specific categories with the most robust associations.
Table 11: Correlation coefficients of linguistic features with the BERT score.
Variable Pearson Spearman
Word count -0.1961* -0.2650*
Polarity -0.1036* -0.1473*
Subjectivity -0.1753* -0.1866*
Readability 0.1182* 0.1518*
* p-value less than 0.01.
Finally, Table 11 examines the potential relationship between the BERT score and various linguistic features automatically extracted from the text. We find statistically significant correlations between the BERT score and all linguistic features analyzed. Notably, there is a strong inverse correlation with both word count and subjectivity, indicating that shorter and more objective texts tend to have higher BERT scores. This finding, together with the correlations presented in Table 9, suggests that BERT is more closely related to linguistic features than to the numerical variables associated with loan applications.
7 Results of the LLM-enhanced granting model
7.1 Analysis of the classification performance
First, we evaluate the impact of incorporating the BERT score in the granting model. In our baseline experiment, we optimize XGBoost with a genetic algorithm using the quantitative and categorical variables typically used in granting models, while in the competing approach, we optimize it but include the BERT score as input variable. Table 12 shows the results of both approaches.
A closer examination of the balanced metrics reveals a marginal enhancement in both BACC (0.6154 vs. 0.6187) and AUC (0.6575 vs. 0.6644), the latter significant according to the DeLong test [46] at the 0.01 level. It is crucial to note that the Lending Club dataset exclusively consists of approved loans. This fact poses a substantial challenge to significantly enhance the outcomes in a loan granting model such as ours since the loans included in the dataset were initially considered favorable by the platform. Furthermore, in experiments not reported in the paper we used CatBoost [47] instead of XGBoost and obtained a similar BACC improvement.
Table 12: Classifier performance with and without the BERT score.
Metric Quant. + Categ. var. Quant. + Categ. + BERT score
BACC 0.6154 0.6187
AUC 0.6575 0.6644
F1 0.3266 0.3308
Precision 0.2168 0.2249
Recall 0.6614 0.6360
Accuracy 0.5835 0.6066
The performance metrics in Table 12 indicate that incorporating the BERT score results in improved precision but diminished recall. However, attributing this change in precision-recall behavior solely to the BERT score requires careful consideration. This caution arises from our observations within our dataset, where classifiers with different hyperparameters and similar near-optimal balanced accuracy values have demonstrated varying precision-recall behaviors, suggesting that this may also be the case here.
Table 13 shows an additional experiment in which XGBoost classifiers are trained and optimized using only one kind of input variable. While the classifier using the quantitative variables is clearly the best, the classifier that uses just the BERT score obtains slightly better results than the one using the four categorical variables (the AUC difference is significant at the 0.01 level according to the DeLong test). This result is noteworthy given the well-known effectiveness and the meaningful nature of the qualitative variables.
Table 13 also shows the result of an XGBoost classifier using the textual features presented in Table 11, namely: polarity, subjectivity, word count, and readability score. This classifier is outperformed by the XGBoost that uses the BERT score in terms of balanced accuracy and AUC (significant at the 0.01 level according to the DeLong test). This finding underscores the superior ability of the fine-tuned LLM to leverage textual descriptions and extract relevant information for the classification task.
Table 13: Performance of the XGBoost considering only a subset of variables.
Metric Quant. Categ. Text. BERT score
BACC 0.6062 0.5486 0.5258 0.5490
AUC 0.6457 0.5656 0.5309 0.5714
F1 0.3192 0.2759 0.2534 0.2601
Precision 0.2138 0.1746 0.1665 0.1877
Recall 0.6300 0.6563 0.5302 0.5153
Accuracy 0.5896 0.4738 0.5227 0.5724
7.2 Feature importance and explainability
To further analyze the impact of the BERT score, in this section, we analyze whether the models’ use of the input variables changes. Figure 4 shows the feature importance given by XGBoost both with and without incorporating the BERT score. Notably, the BERT score emerges as the third most influential variable, accounting for 6% importance. Moreover, the order and magnitude of importance for other variables exhibit shifts. Specifically, the borrower’s annual revenue diminishes in importance following the inclusion of the BERT score, while the significance of the ‘credit card’ purpose increases. This observation underscores the significant influence of the BERT score on the classifier.
(a) Before BERT score incorporation.
(b) After BERT score incorporation.
Figure 4: Feature importances of the XGBoost classifiers.
To gain a deeper understanding of the role of variables, we employ SHAP values [48]. Figure 5 provides a visualization of the SHAP values for the 10 variables with the most significant impact on the model’s output, comparing models with and without including the BERT score.
(a) Before BERT score incorporation.
(b) After BERT score incorporation.
Figure 5: SHAP values of the XGBoost classifiers.
It is noteworthy that Figure 5(a) resembles a comparable figure presented in [41], wherein SHAP values were utilized in a lending model on Lending Club [^11] Our dataset is more constrained, only considering loans with textual descriptions, as explained in Section 4.. Examining Figure 5(b), we find that the BERT score ranks as the fourth variable in terms of impact. A closer inspection of the violin plots reveals that including BERT does not substantially alter how variables influence the target class.
Figure 6: Dependence plot between SHAP values and BERT score.
Figure 6 reveals a direct and linear relationship between the BERT score and the default risk. Notably, this relationship is asymmetric around BERT score values of 0.5; scores below 0.4 correspond to SHAP values ranging between -0.4 and -0.6, leading the model to predict non-default. Conversely, this impact range in the positive case is only reached by BERT scores exceeding 0.7, signifying that only exceptionally high BERT scores serve as strong indicators of default.
7.3 Impact across the purpose categories
Now we delve into the influence of incorporating the BERT score on classification outcomes across various categories of the purpose variable. Leveraging the language comprehension capabilities of LLMs, it is conceivable that the BERT model can provide more accurate characterizations of loan purposes than the purpose variable alone. For instance, the ‘other’ category, being inherently ambiguous, could benefit from a more nuanced understanding of the loan description, potentially leading to improved prediction outcomes.
Figure 7: BACC changes by purpose after including the BERT score.
Figure 7 illustrates the changes in balanced accuracy upon incorporating the BERT score for each loan purpose. Notably, the ‘other’ category shows a modest improvement of 2.11%. However, several other categories exhibit more substantial enhancements, including ‘educational’ (9.22%), ‘moving’ (5.84%), ‘medical’ (3.84%), and ‘small business’ (3.40%). Given the black-box nature of our model, it is difficult to ascertain the reasons why these categories have improved more than the others. However, we posit that these categories share a commonality—the more detailed specification of purpose or a deeper understanding of the borrower’s situation contributes to a more precise delineation of the default risk. For instance, in education loans, factors like the field of study and the educational institution may play a crucial role in determining employability and salaries, thereby enabling a more accurate assessment of the default risk. The same happens in the case of moving to a new place, having to pay medical expenses, or obtaining funds for a business.
The following examples illustrate instances of loans that were initially misclassified as defaults in the absence of the BERT score but were subsequently correctly predicted as non-defaults after its inclusion:
Purpose: Educational; BERT score: 0.3765
“I’m 25 years old and living in New Orleans. I’m asking for a relatively small amount of money to help me take care of my post-Bac tuition for teacher certification and to help me pay off a credit card. I currently work as a private school teacher making very little money with no benefits (about \$29,000 a year). I have to pay about \$1000 in the coming year for my tuition, and I have to get health insurance ASAP, but it’s hard to do so with no financial help from anyone else. My parents can’t help me because my mother is permanently disabled and my father took a huge pay cut this year.”
Purpose: Moving; BERT score: 0.3843
“Although I can afford payments,due to some recent expenses, I am short on cash flow for an unexpected move. I am, however, looking for a more reasonable alternative to banking rates. I have borrowed from Lending Club before and always paid fully and on time with automatic payments.”
Purpose: Small business; BERT score: 0.3292
“The purpose of this loan is to fund advertising costs for a growing internet business venture. I am a successful sales professional earning an average of over \$250K per year over the last 5 years. My credit scores are strong and I have a documented history of paying all my debts (personal or business related) on time.”
Purpose: Medical; BERT score: 0.3921
“This loan will be used to pay off a Care Credit credit card currently at 21.9% I used the card to pay for a prosthetic limb that my health insurer would not cover.”
In each of these examples, the corresponding BERT scores fall below 0.4. As demonstrated in Figure 6, the classifier will have a stronger driver to categorize the loan as non-default when the score is below this value. As all the examples offer more details on the situation and the purpose of the loan, their further clarity could explain their low BERT score. In addition, all the texts shown are accurate, readable, and written in a confident and objective tone. While previous evidence in the literature offered mixed results about the influence of linguistic factors as default predictors [3, 18, 4, 7], we have shown that our BERT score is significantly related to them. While these explanations are sound, given the black-box nature of our BERT-based approach, the explanations influencing the BERT score of each text cannot be precisely determined. This aspect needs to be improved to comply with the requirements of the regulators and foster trust among the end-users.
8 Conclusions
In this paper, we have introduced a novel approach that highlights the value of state-of-the-art natural language processing techniques in enhancing credit risk models. Our approach fine-tunes BERT using loan descriptions to generate a risk score that effectively discriminates between defaulted and non-defaulted loans. Furthermore, we show that integrating the BERT-based risk score with traditional variables significantly enhances the performance of conventional granting models. These results underscore the potential of LLMs to augment lending models, emphasizing their invaluable role in harnessing the richness of textual data to improve credit risk assessment.
Our approach can be easily applied without manual annotation, which is a complex task that also introduces annotator subjectivity. As our analysis shows, the score incorporates both linguistic factors and content aspects that relate to the creditworthiness of the borrower’s situation and purpose. Additionally, while fine-tuning the LLM is a time-consuming process that requires GPU processing, generating predictions, such as the risk score for a loan description, can be done instantly with standard home equipment.
Our work represents just the first steps in this endeavor and underscores the necessity for further exploration and integration of cutting-edge methodologies to enhance the effectiveness of risk-return assessments in the dynamic landscape of P2P lending or even in traditional financial institutions. Several directions deserve to be explored to refine the predictive capability of the model and the understanding of loan applicants’ conditions.
An important limitation of our work is the lack of transparency regarding the factors contributing to a given risk score and its potential biases. These issues hinder its application in real-world contexts, where a comprehensive understanding of the score generation process and assurance that it does not introduce biases related to gender, ethnicity, financial inclusion, or other factors are essential. Thus, efforts to enhance the transparency and explainability of the risk score are crucial not only for regulatory compliance but also for fostering trust among borrowers and lenders [49].
At this respect, LLM-based topic modeling tools such as BERTopic [50] could be used to gain an understanding of the risk scores. For example, topic modeling could identify whether a loan description is related to “risky topics” that in turn could be incorporated as input variables in the granting model. Similarly, the embeddings from BERT or other LLMs could be used to generate interpretable topics from large, complex vocabularies, as other works have done [51]. Applying such topic modeling approaches to loan descriptions could facilitate the identification of situations with varying risk levels. These topics could then serve as additional variables in classification models, improving predictive performance while maintaining the transparency required in financial assessments.
Another potential research avenue involves delving into more advanced LLMs. For instance, encoder-only models like RoBERTa [37] could improve the model’s capacity to capture more intricate linguistic patterns. Furthermore, the rise of generative AI has spurred interest in exploring decoder models for classification purposes. Although these models were not originally designed for such tasks, recent advancements have shown promising approximations. For example, CARP (Clue And Reasoning Prompting) [52] utilizes in-context learning, generating textual responses conditioned on a provided prompt with a few annotated examples (few-shot learning). Remarkably, this approach improved results in classification tasks without the need for transfer learning and fine-tuning and could be applied to risk scoring.
Future work should also address the economic implications of our findings by integrating cost- or profit-sensitive approaches [53, 42]. These methods allow for the inclusion of the financial impacts of classification outcomes. Investigating whether improvements in model performance by including textual descriptions translate into better financial outcomes would enhance the real-world applicability of our findings.